Al the beginning of the year, Swit, Incs Work in Process Inventory account had a balance of $120.000 t n production, and $73000 of direct labor costs were incurred Manufacturing overhead amounted to $828000 The cost of goods manulachured was $673,000 Wh is the balance in the Work in Process Inventory account on December 31 uring the year, $250,000 of drect materialh wore ned O A. 5673.000 OB. 5590,000 O C. $1,346,000 OD. $1.151,000 ect

Viva, Inc. has provided the following information for the year Cost of Goods Manufactured Beginning Balance Finished Goods Inventory98,000 Ending Balance Finished Goods Inventory $1,249,000 82,000 What is the cost of goods sold? O A. $1,233,000 OB. $1,249,000 O C. $1,265,000 D. $180,000 ct

herch Five Seasons is a merchandiser of packed foods. The company provides the following information for the year Reven $160,000 61,000 68,000 31,000 28,000 ber of Sales Revenue Cost of Goods Sold Operating Expenses Net Income Number of Units Sold How much was the unit cost per item of product sold? (Round your answer to the nearest cent) Pi id is A $10111 ibution B. $461 gil oc. s571 co D $2 18 sells ixed come f
Answer #1

Answer 1:

Correct answer is:

B. $598,000

Explanation:

Ending Work in process = Direct material + Direct labor + Manufacturing overhead + Beginning Work in process - Cost of goods manufactured

= 250000 + 73000 + 828000 + 120000 - 673000

=$598,000

As such option B is correct and options A, C and D are incorrect.

Answer 2:

Correct answer is:

C. $1,265,000

Explanation:

Cost of goods sold = Cost of goods manufactured + Beginning balance Finished goods inventory - Ending Balance Finished goods inventory

= 1249000 + 98000 - 82000

= $1,265,000

As such option C is correct and options A, B and D are incorrect.

Answer 3:

Correct answer is:

D. $2.18

Explanation:

Unit cost of per item of product sold = Cost of goods sold / Number of units sold

= 61000 / 28000

= $2.18

As such option D is correct and options A, B and C are incorrect.